# Dark matter cooling during early matter-domination boosts sub-earth halos

### v1

@@ Abstract
The existence of an early matter-dominated epoch prior to the big bang nucleosynthesis may lead to a scenario where the thermal dark matter cools faster than plasma before the radiation dominated era begins. In the radiation-dominated epoch, dark matter free-streams after it decouples both chemically and kinetically from the plasma. In the presence of an early matter-dominated era, chemical decoupling of the dark matter may succeed by a partial kinetic decoupling before reheating ends, depending upon the contributions of different partial wave amplitudes in the elastic scattering rate of the dark matter. We show that the $s$-wave scattering is sufficient to partially decouple the dark matter from the plasma, if the entropy injection during the reheating era depends on the bath temperature, while $p$-wave scattering leads to full decoupling in such cosmological backdrop. The decoupling of dark matter before the end of reheating causes an additional amount of cooling, reducing its free-streaming horizon compared to usual radiation-dominated cosmology. The enhanced matter perturbations for scales entering the horizon prior to the end of reheating, combined with the reduced free-steaming horizon, increase the number density of sub-earth mass halos. Resulting boost in the dark matter annihilation signatures could offer an intriguing probe to differentiate pre-BBN non-standard cosmological epochs. We show that the free-streaming horizon of the dark matter requires to be smaller than a cut-off to ensure boost in the sub-earth halo populations. As case studies we present two examples: one for a scalar dark matter with $s$-wave elastic scattering and the other one featuring a fermionic dark matter with $p$-wave elastic scattering. We identify regions of parameter space in both models where the dark matter kinetically decouples during reheating, amplifying small scale structure formation.
@@ 1 Introduction
Testability of weakly interacting massive particles (WIMPs) through a variety of direct and indirect detection experiments establish them as one of the most promising candidates for dark matter [15, 16, 17, 18, 19], in the mass range between a few GeV to TeV. In the WIMP scenario, DM remains in equilibrium with the thermal plasma of visible matter in the early universe. As the universe expands, after some time the interaction rate ($\Gamma$) of DM with particles in the thermal bath becomes ineffective (i.e., $\Gamma<H$, where $H$ is the Hubble expansion rate) to keep it in thermal equilibrium with the plasma. As a result the DM thermally decouples from the standard model plasma. The thermal decoupling results in a freeze-out of the DM, and it sets the relic density of DM, which at the present day has a value of $\Omega_{\rm DM}^{0}h^{2}\simeq 0.12$, as measured by the Planck collaboration [2].
@@
In the WIMP scenario, it is important to note that the thermal decoupling of DM implies both chemical and kinetic decoupling from the plasma. Chemical decoupling of DM sets the number density of relic DM in the universe, however, full thermal decoupling of DM is crucial for forming structures. While chemical decoupling is primarily governed by the annihilation of the DM ($\chi$) into the bath particles ($B$), represented as $\chi\chi\to BB$, the kinetic decoupling depends on the elastic scattering processes, i.e., $\chi B\to\chi B$. In general, chemical and kinetic decoupling may not take place at the same time, since the annihilation rate of the DM $(\Gamma_{\rm ann}=n^{\rm eq}_{\chi}\langle\sigma v\rangle_{\rm ann})$ and the elastic scattering rate $(\Gamma_{\rm el}=n^{\rm eq}_{B}\langle\sigma v\rangle_{\rm el})$ vary with the bath temperature in a qualitatively different manner. For a cold dark matter, i.e., if the freeze-out is non-relativistic, the equilibrium number density of the DM after the freeze-out falls as $n^{\rm eq}_{\chi}\sim T^{3/2}e^{-m_{\chi}/T}$, where $m_{\chi}$ is the mass of the DM and $T$ represents the temperature of the bath. In contrast, the relativistic bath particles’ number density varies as $n^{\rm eq}_{B}\sim T^{3}$. Thus, in a radiation dominated universe ($H\sim T^{2}/M_{\rm pl}$), chemical decoupling precedes the kinetic decoupling of cold dark matter, assuming that the thermal averaged cross-section mildly varies with temperature.
@@
In the standard lore, the DM is presumed to undergo decoupling in a universe dominated by radiation prior to the commencement of Big Bang nucleosynthesis (BBN). Nevertheless, hardly any evidence exists to confirm that the universe was only radiation dominated prior to the BBN, since the end of inflation. This ambiguity prompts various conjectures regarding the alternative thermal evolution of the cosmos, a pre-BBN early matter-dominated epoch (EMDE) emerging as a prominent hypothesis. Several motivated scenarios beyond the Standard Model (BSM) propose the existence of meta-stable oscillating scalar fields, such as string theory inspired moduli fields [20, 21, 22], supersymmetric condensates and gravitino [23, 24], scenarios with primordial black hole formation [25, 26, 27], inflaton fields [28, 29], curvaton [30], dilaton [31], and Q-balls [32], among others. These long-lived species could instigate an early matter-dominated epoch after inflation, followed by reheating due to the late decay of these oscillating matter fields into radiation [33, 34, 35, 36, 37, 38]. The presence of such an EMDE strongly impacts both the thermal and non-thermal DM production [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77], and their decoupling [78, 79, 80].
@@
In such an EMDE scenario, the $s$-wave elastic scattering of non-relativistic DM ($\langle\sigma v\rangle_{\rm el}\sim{\rm const.}$) is insufficient to kinetically decouple the DM from the plasma before the onset of radiation dominated epoch, as the momentum exchange rate in elastic scattering ($\gamma_{\rm el}=\Gamma_{\rm el}T/m_{\chi}$) and the Hubble expansion both scale as $T^{4}$. Thus, the scales entering the horizon prior to the end of reheating can not lead to the formation of protohalos [81, 82, 83]. However, $p$-wave elastic scattering ($\langle\sigma v\rangle_{\rm el}\sim T^{2}$) yielding $\gamma_{\rm el}\sim T^{6}$ leads to a partial kinetic decoupling, as shown in [80]. Consequently, the DM begins to cool at a faster rate than the plasma during the EMDE. Eventually, after reheating, DM fully decouples from the plasma. Due to this partial decoupling, the additional cooling of DM during EMDE reduces the free-streaming horizon compared to the standard scenario of DM decoupling in the radiation-dominated epoch. Thus, the free-streaming of DM in this type of non-standard scenario does not suppress small-scale structures (sub-earth halos) formed by the scales entering the horizon before reheating. On the contrary, such small scales are additionally endowed with enhanced matter perturbations in the presence of an EMDE causing the formation of enhanced structures compared to the standard scenario [84, 85, 86].
@@
As a proof of principle, we consider two simple models of dark matter, one with a scalar DM and the other with a fermionic DM, where the DM elastically scatters off radiation through $s$-wave and $p$-wave partial wave amplitude channels, respectively. For both cases, we demonstrate that the parameter space that satisfies the observed relic density can also lead to kinetic decoupling of the DM during the EMDE yielding a reduced free-streaming horizon of the DM compared to purely radiation dominated cosmology. Additionally, we demonstrate that, despite kinetic decoupling during the EMDE, the enhancement in small-scale structure formation occurs only within a specific region of parameter space where the free-streaming horizon falls below a certain cut-off value.
The formation of microhalos may lead to an enhanced annihilation rate of DM, providing an intriguing probe for pre-BBN thermal history. The detection of gamma rays from DM annihilation at the galactic centre by the indirect detection experiments [88, 89, 90], such as the Fermi-LAT [91, 92, 93] place strong constraints on the DM annihilation rate. In the presence of an EMDE, two compensatory effects determine the overall annihilation rate of DM: (i) a reduced annihilation cross-section of thermal DM to match the correct relic density due to entropy dilution, and (ii) a boost in the number density of microhalos resulting from enhanced matter perturbations and additional cooling of partially decoupled DM before reheating. While we show the enhancement of microhalo population due to the presence of an EMDE with temperature dependent entropy injection, a full dedicated study to compute the boost factor for DM annihilation rate signatures is deferred to a future work [94]. Besides annihilation signatures of DM [95, 96, 97, 98], sub-halo collisions with old neutron stars, sub-halo dark matter scatterings with cosmic rays [99],microlensing [100, 101], and gravitational waves produced in EMDE [102, 103, 104, 105, 106, 107, 108] could offer alternative methods to investigate non-standard cosmological epochs, due to an increased sub-earth halo populations, as described in this paper.
The paper is organized as follows. In Sec. 2 we discuss the mechanism of kinetic decoupling in different cosmological backdrops and compute the free-streaming horizon of the DM. In Sec. 3, we first describe the thermal history of an EMDE with temperature dependent entropy injection. The rest of the section is dedicated to a detailed study of the cosmological perturbations during the EMDE and formation of sub-earth mass halos. Two simple models featuring a scalar and a fermionic dark matter which exhibit $s$-wave and $p$-wave elastic scatterings, respectively, and decoupling during the EMDE are explored as case studies in Sec. 4. We conclude in Sec. 5.
@@ 2 Partial kinetic decoupling of dark matter
where $\mathcal{C}_{\rm ann}[f_{\chi}]$ and $\mathcal{C}_{\rm el}[f_{\chi}]$ denote the collision terms corresponding to the annihilation ($\chi\chi\to BB$) and the elastic scatterings ($\chi B\to\chi B$) of the DM with the bath particles ($B$), respectively. The number density ($n_{\chi}$) and the temperature ($T_{\chi}$) of the DM are defined by taking the moments of the phase space distribution as [111, 79, 112]
@@
where $K_{1,2}$ are the modified Bessel function of the second kind. Once the universe cools down to $T_{\rm fo}$, DM leaves the chemical equilibrium and the co-moving number density of the DM freezes-out.
@@
To arrive at this equation, we specifically assume that the DM is non-relativistic below the freeze-out temperature ($m_{\chi}>T_{\rm fo}$). As long as the DM is thermally coupled to the plasma, $T_{\chi}$ is equal to the temperature of the plasma ($T$). When the elastic scattering rate becomes smaller than the Hubble expansion rate i.e., $\gamma_{\rm el}(T)\ll H(T)$, the solution of Eq. (2.7) shows that $T_{\chi}$ redshifts as $T_{\chi}\sim a^{-2}$. This solution additionally necessitates another condition, i.e., $\gamma_{\rm el}(T)T\ll HT_{\chi}$. We will explore the implications of relaxing this condition in the subsequent sections. In a radiation-dominated universe, the bath temperature evolves as $T\sim a^{-1}$ while the DM temperature cools as $T_{\chi}\sim a^{-2}$ after kinetic decoupling. However, the situation becomes more involved if the thermal history of the universe during the decoupling is different from the usual radiation domination [78, 79, 80]. In the following we investigate the kinetic decoupling of the DM for non-standard cosmological backdrops.
@@ 2.1 Kinetic decoupling in non-standard cosmological scenario
We define $T_{\rm dec}$ ($a_{\rm dec}$) as the temperature (scale factor) when $\gamma_{\rm el}(T_{\rm dec})=H(T_{\rm dec})$. As indicated earlier, the condition for kinetic decoupling, $\gamma_{\rm el}(T)\ll H$, does not necessarily imply that the right hand side of Eq. (2.7) is also negligible compared to $T_{\chi}$. In fact, the approximate solution of Eq. (2.7) in the limit $\gamma_{\rm el}(T)\ll H$, but $\gamma_{\rm el}(T)T\nless HT_{\chi}$ is given by
@@
Approximate solution of these equations yield
@@
Table 1: Conditions for kinetic decoupling of the DM are shown for different non-standard cosmological scenarios with entropy injection. For the purpose of illustration some benchmark values of $k$ and $m$ are chosen for three different values of $\omega_{\phi}=0,1/3,1$, covering various possibilities of kinetic decoupling. In the last two columns we specifically consider the cases where $s$-wave ($n=0$) and $p$-wave ($n=2$) scatterings are dominant.
@@ 2.2 Free-streaming of partially decoupled dark matter
Once chemically and kinetically decoupled from the thermal bath, the dark matter starts to free-stream with a velocity $v_{\chi}(a)\propto\sqrt{T_{\chi}(a)}$. The free-streaming horizon $\lambda_{\rm fsh}$ of the DM sets the scale below which the structures will be washed out due to the large DM velocity. The scaling of $T_{\chi}(a)$ with the scale factor($a$) plays a crucial role to determine $\lambda_{\rm fsh}$. We compute $\lambda_{\rm fsh}$ in two cases, namely, in presence of an EMDE, and in purely radiation dominated (RD) universe using the general solution Eq. (2.9) for $T_{\chi}(a)$ and the following expressions
@@
We define $T_{\rm kds}$ as the temperature at kinetic decoupling in the radiation dominated epoch, such that $\gamma_{\rm el}(T_{\rm kds})=H_{\rm RD}(T_{\rm kds})$, while $T_{\rm dec}$ corresponds to the decoupling temperature in the EMDE when $\gamma_{\rm el}(T_{\rm dec})=H_{\rm EMD}(T_{\rm dec})$.
@@
We observe that in all three cases, the free-streaming horizon is smaller in the case of EMDE if $\gamma_{0}^{\rm EMD}\leq\gamma_{0}^{\rm RD}$, (i.e. $r\leq 1$). If the primary interaction governing the relic density of the DM also maintains it in kinetic equilibrium with the bath, one typically needs $r<1$ to match the observed relic density in both the early matter-dominated and the radiation-dominated cosmology. This requirement stems from the fact that the entropy injection during the EMDE phase dilutes the relic density, necessitating a comparatively lower interaction rate. However, in the presence multiple interactions, one may fathom a situation where the processes involved in setting the relic density may not be the dominant interactions required for maintaining the kinetic equilibrium, such that $r>1$ is possible. In that case, $\lambda^{\rm EMD}_{\rm fsh}$ is smaller than $\lambda^{\rm RD}_{\rm fsh}$, only if $T_{\rm dec}/T_{\rm RH}$ is greater than a certain threshold value.
Figure 1: The ratios of dark matter free-streaming horizon in presence of an EMDE and in standard radiation dominated epoch are shown in $T_{\rm kds}-T_{\rm dec}$ plane in the units of $T_{\rm RH}$. We fix the value of $T_{\rm RH}$ at 10 MeV for illustration. The red lines represent the isocontours for the ratio $r$. For $r<1$, the free-streaming horizon in an EMDE is always smaller than its counterpart in RD era. The change in the slope of the isocontours of $r$ around $T_{\rm dec}/T_{\rm RH}\sim 20$ is due to the jump in $g_{*}(T)$ around the QCD phase transition.
In Sec. 4, we will consider two simple models of dark matter where the DM scatters off the bath particles through $s$-wave and $p$-wave scattering, and compute the free-streaming horizon in terms of microscopic model parameters, in the presence of an EMDE with time dependent entropy injection.

### v2

@@ Abstract
The existence of an early matter-dominated epoch prior to the Big Bang Nucleosynthesis (BBN) may lead to a scenario where the thermal dark matter cools faster than plasma before the radiation-dominated era begins. In the radiation-dominated epoch, dark matter free-streams after it decouples both chemically and kinetically from the plasma. In the presence of an early matter-dominated era, chemical decoupling of the dark matter may succeed by a partial kinetic decoupling before reheating ends, depending upon the contributions of different partial wave amplitudes in the elastic scattering rate of the dark matter. We show that the $s$-wave scattering is sufficient to partially decouple the dark matter from the plasma, if the entropy injection during the reheating era depends on the bath temperature, while $p$-wave scattering leads to full decoupling in such cosmological backdrop. The decoupling of dark matter before the end of reheating causes an additional amount of cooling, reducing its free-streaming horizon compared to the usual radiation-dominated cosmology. The enhanced matter perturbations for scales entering the horizon prior to the end of reheating, combined with the reduced free-steaming horizon, increase the number density of sub-earth mass halos. The resulting boost in the dark matter annihilation signatures could offer an intriguing probe to differentiate pre-BBN non-standard cosmological epochs. We show that the free-streaming horizon of the dark matter requires to be smaller than a cut-off to ensure a boost in the sub-earth halo populations. As case studies, we present two examples: one for a scalar dark matter with $s$-wave elastic scattering and the other one featuring a fermionic dark matter with $p$-wave elastic scattering. We identify regions of parameter space in both models where the dark matter kinetically decouples during reheating, amplifying small-scale structure formation.
@@ 1 Introduction
Testability of weakly interacting massive particles (WIMPs) through a variety of direct and indirect detection experiments establish them as one of the most promising candidates for dark matter [15, 16, 17, 18, 19], in the mass range between a few GeV to TeV. In the WIMP scenario, the DM remains in equilibrium with the thermal plasma of visible matter in the early universe. As the universe expands, after some time the interaction rate ($\Gamma$) of DM with particles in the thermal bath becomes ineffective (i.e., $\Gamma<H$, where $H$ is the Hubble expansion rate) to keep it in thermal equilibrium with the plasma. As a result, the DM thermally decouples from the standard model plasma. The thermal decoupling results in a freeze-out of the DM, and it sets the relic density of DM, which at the present day has a value of $\Omega_{\rm DM}^{0}h^{2}\simeq 0.12$, as measured by the Planck collaboration [2].
@@
In the WIMP scenario, it is important to note that the thermal decoupling of DM implies both chemical and kinetic decoupling from the plasma. Chemical decoupling of DM sets the number density of relic DM in the universe, however, full thermal decoupling of DM is crucial for forming structures. While chemical decoupling is primarily governed by the annihilation of the DM ($\chi$) into the bath particles ($B$), represented as $\chi\chi\to BB$, the kinetic decoupling depends on the elastic scattering processes, i.e. $\chi B\to\chi B$. In general, chemical and kinetic decoupling may not take place at the same time, since the annihilation rate of the DM $(\Gamma_{\rm ann}=n^{\rm eq}_{\chi}\langle\sigma v\rangle_{\rm ann})$ and the elastic scattering rate $(\Gamma_{\rm el}=n^{\rm eq}_{B}\langle\sigma v\rangle_{\rm el})$ vary with the bath temperature in a qualitatively different manner. For a cold dark matter, i.e., if the freeze-out is non-relativistic, the equilibrium number density of the DM after the freeze-out falls as $n^{\rm eq}_{\chi}\sim T^{3/2}e^{-m_{\chi}/T}$, where $m_{\chi}$ is the mass of the DM and $T$ represents the temperature of the bath. In contrast, the relativistic bath particles’ number density varies as $n^{\rm eq}_{B}\sim T^{3}$. Thus, in a radiation-dominated universe ($H\sim T^{2}/M_{\rm pl}$), chemical decoupling precedes the kinetic decoupling of cold dark matter, assuming that the thermal averaged cross-section mildly varies with temperature.
@@
In the standard lore, the DM is presumed to undergo decoupling in a universe dominated by radiation prior to the commencement of Big Bang Nucleosynthesis (BBN). Nevertheless, hardly any evidence exists to confirm that the universe was only radiation-dominated prior to the BBN, since the end of inflation. This ambiguity prompts various conjectures regarding the alternative thermal evolution of the cosmos, a pre-BBN early matter-dominated epoch (EMDE) emerging as a prominent hypothesis. Several motivated scenarios beyond the Standard Model (BSM) propose the existence of meta-stable oscillating scalar fields, such as string theory inspired moduli fields [20, 21, 22], supersymmetric condensates and gravitino [23, 24], scenarios with primordial black hole formation [25, 26, 27], inflaton fields [28, 29], curvaton [30], dilaton [31], and Q-balls [32], among others. These long-lived species could instigate an early matter-dominated epoch after inflation, followed by reheating due to the late decay of these oscillating matter fields into radiation [33, 34, 35, 36, 37, 38]. The presence of such an EMDE strongly impacts both the thermal and non-thermal DM production [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] and their decoupling [79, 80, 81].
@@
In such an EMDE scenario, the $s$-wave elastic scattering of non-relativistic DM ($\langle\sigma v\rangle_{\rm el}\sim{\rm const.}$) is insufficient to kinetically decouple the DM from the plasma before the onset of radiation-dominated epoch, as the momentum exchange rate in elastic scattering ($\gamma_{\rm el}=\Gamma_{\rm el}T/m_{\chi}$) and the Hubble expansion both scale as $T^{4}$. Thus, the scales entering the horizon prior to the end of reheating can not lead to the formation of protohalos [82, 83, 84]. However, $p$-wave elastic scattering ($\langle\sigma v\rangle_{\rm el}\sim T^{2}$) yielding $\gamma_{\rm el}\sim T^{6}$ leads to a partial kinetic decoupling, as shown in [81]. Consequently, the DM begins to cool at a faster rate than the plasma during the EMDE. Eventually, after reheating, DM fully decouples from the plasma. Due to this partial decoupling, the additional cooling of DM during EMDE reduces the free-streaming horizon compared to the standard scenario of DM decoupling in the radiation-dominated epoch. Thus, the free-streaming of DM in this type of non-standard scenario does not suppress small-scale structures (sub-earth halos) formed by the scales entering the horizon before reheating. On the contrary, such small scales are additionally endowed with enhanced matter perturbations in the presence of an EMDE causing the formation of enhanced structures compared to the standard scenario [85, 86, 87].
@@
As a proof of principle, we consider two simple models of dark matter, one with a scalar DM and the other with a fermionic DM, where the DM elastically scatters off radiation through $s$-wave and $p$-wave partial wave amplitude channels, respectively. For both cases, we demonstrate that the parameter space that satisfies the observed relic density can also lead to kinetic decoupling of the DM during the EMDE yielding a reduced free-streaming horizon of the DM compared to purely radiation-dominated cosmology. Additionally, we demonstrate that, despite kinetic decoupling during the EMDE, the enhancement in small-scale structure formation occurs only within a specific region of parameter space where the free-streaming horizon falls below a certain cut-off value.
The formation of microhalos may lead to an enhanced annihilation rate of DM, providing an intriguing probe for pre-BBN thermal history. The detection of gamma rays from DM annihilation at the galactic center by the indirect detection experiments [89, 90, 91], such as the Fermi-LAT [92, 93, 94] place strong constraints on the DM annihilation rate. Ref. [95] further places a constraint on the reheating temperature from the WIMP dark matter annihilation using the Fermi-LAT data, assuming a fixed mass for the microhalos. In the presence of an EMDE, two compensatory effects determine the overall annihilation rate of DM: (i) a reduced annihilation cross-section of thermal DM to match the correct relic density due to entropy dilution, and (ii) a boost in the number density of microhalos resulting from enhanced matter perturbations and additional cooling of partially decoupled DM before reheating. While we show the enhancement of microhalo population due to the presence of an EMDE with temperature-dependent entropy injection, a full dedicated study to compute the boost factor for DM annihilation rate signatures is deferred to a future work [96]. Besides annihilation signatures of DM [97, 98, 99, 100, 101], sub-halo collisions with old neutron stars, sub-halo dark matter scatterings with cosmic rays [102], pulsar timing arrays [103, 104, 95, 105], microlensing [106, 107], and gravitational waves produced in EMDE [108, 109, 110, 111, 112, 113, 114] could offer alternative methods to investigate non-standard cosmological epochs, due to an increased sub-earth halo population, as described in this paper.
The paper is organized as follows. In Sec. 2 we discuss the mechanism of kinetic decoupling in different cosmological backdrops and compute the free-streaming horizon of the DM. In Sec. 3, we first describe the thermal history of an EMDE with temperature-dependent entropy injection. The rest of the section is dedicated to a detailed study of the cosmological perturbations during the EMDE and the formation of sub-earth mass halos. Two simple models featuring a scalar and a fermionic dark matter which exhibit $s$-wave and $p$-wave elastic scatterings, respectively, and decoupling during the EMDE are explored as case studies in Sec. 4. We conclude in Sec. 5.
@@ 2 Partial kinetic decoupling of dark matter
where $\mathcal{C}_{\rm ann}[f_{\chi}]$ and $\mathcal{C}_{\rm el}[f_{\chi}]$ denote the collision terms corresponding to the annihilation ($\chi\chi\to BB$) and the elastic scatterings ($\chi B\to\chi B$) of the DM with the bath particles ($B$), respectively. The number density ($n_{\chi}$) and the temperature ($T_{\chi}$) of the DM is defined by taking the moments of the phase space distribution as [117, 80, 118]
@@
where $K_{1,2}$ are the modified Bessel function of the second kind. Once the universe cools down to $T_{\rm fo}$, DM leaves the chemical equilibrium, and the co-moving number density of the DM freezes out.
@@
To arrive at this equation, we specifically assume that the DM is non-relativistic below the freeze-out temperature ($m_{\chi}>T_{\rm fo}$). As long as the DM is thermally coupled to the plasma, $T_{\chi}$ is equal to the temperature of the plasma ($T$). When the elastic scattering rate becomes smaller than the Hubble expansion rate i.e., $\gamma_{\rm el}(T)\ll H(T)$, the solution of Eq. (2.7) shows that $T_{\chi}$ redshifts as $T_{\chi}\sim a^{-2}$. This solution additionally necessitates another condition, i.e., $\gamma_{\rm el}(T)T\ll HT_{\chi}$. We will explore the implications of relaxing this condition in the subsequent sections. In a radiation-dominated universe, the bath temperature evolves as $T\sim a^{-1}$ while the DM temperature cools as $T_{\chi}\sim a^{-2}$ after kinetic decoupling. However, the situation becomes more involved if the thermal history of the universe during the decoupling is different from the usual radiation domination [79, 80, 81]. In the following, we investigate the kinetic decoupling of the DM for non-standard cosmological backdrops.
@@ 2.1 Kinetic decoupling in non-standard cosmological scenario
We define $T_{\rm dec}$ ($a_{\rm dec}$) as the temperature (scale factor) when $\gamma_{\rm el}(T_{\rm dec})=H(T_{\rm dec})$. As indicated earlier, the condition for kinetic decoupling, $\gamma_{\rm el}(T)\ll H$, does not necessarily imply that the right-hand side of Eq. (2.7) is also negligible compared to $T_{\chi}$. In fact, the approximate solution of Eq. (2.7) in the limit $\gamma_{\rm el}(T)\ll H$, but $\gamma_{\rm el}(T)T\nless HT_{\chi}$ is given by
@@
The approximate solution of these equations yields
@@
Table 1: Conditions for kinetic decoupling of the DM are shown for different non-standard cosmological scenarios with entropy injection. For the purpose of illustration, some benchmark values of $k$ and $m$ are chosen for three different values of $\omega_{\phi}=0,1/3,1$, covering various possibilities of kinetic decoupling. In the last two columns, we specifically consider the cases where $s$-wave ($n=0$) and $p$-wave ($n=2$) scatterings are dominant.
@@ 2.2 Free-streaming of partially decoupled dark matter
Once chemically and kinetically decoupled from the thermal bath, the dark matter starts to free-stream with a velocity $v_{\chi}(a)\propto\sqrt{T_{\chi}(a)}$. The free-streaming horizon $\lambda_{\rm fsh}$ of the DM sets the scale below which the structures will be washed out due to the large DM velocity. The scaling of $T_{\chi}(a)$ with the scale factor($a$) plays a crucial role in determining $\lambda_{\rm fsh}$. We compute $\lambda_{\rm fsh}$ in two cases, namely, in the presence of an EMDE, and in a purely radiation-dominated (RD) universe using the general solution Eq. (2.9) for $T_{\chi}(a)$ and the following expressions
@@
Figure 1: The ratios of dark matter free-streaming horizon in the presence of an EMDE and in standard radiation dominated epoch are shown in $T_{\rm kds}-T_{\rm dec}$ plane in the units of $T_{\rm RH}$. We fix the value of $T_{\rm RH}$ at 10 MeV for illustration. The red lines represent the isocontours for the ratio $r$. For $r<1$, the free-streaming horizon in an EMDE is always smaller than its counterpart in RD era. The change in the slope of the isocontours of $r$ around $T_{\rm dec}/T_{\rm RH}\sim 20$ is due to the jump in $g_{*}(T)$ around the QCD phase transition.
We define $T_{\rm kds}$ as the temperature at kinetic decoupling in the radiation-dominated epoch, such that $\gamma_{\rm el}(T_{\rm kds})=H_{\rm RD}(T_{\rm kds})$, while $T_{\rm dec}$ corresponds to the decoupling temperature in the EMDE when $\gamma_{\rm el}(T_{\rm dec})=H_{\rm EMD}(T_{\rm dec})$.
@@
We observe that in all three cases, the free-streaming horizon is smaller in the case of EMDE if $\gamma_{0}^{\rm EMD}\leq\gamma_{0}^{\rm RD}$, (i.e., $r\leq 1$). If the primary interaction governing the relic density of the DM also maintains it in kinetic equilibrium with the bath, one typically needs $r<1$ to match the observed relic density in both the early matter-dominated and the radiation-dominated cosmology. This requirement stems from the fact that the entropy injection during the EMDE phase dilutes the relic density, necessitating a comparatively lower interaction rate. However, in the presence of multiple interactions, one may fathom a situation where the processes involved in setting the relic density may not be the dominant interactions required for maintaining the kinetic equilibrium, such that $r>1$ is possible. In that case, $\lambda^{\rm EMD}_{\rm fsh}$ is smaller than $\lambda^{\rm RD}_{\rm fsh}$, only if $T_{\rm dec}/T_{\rm RH}$ is greater than a certain threshold value.
In Sec. 4, we will consider two simple models of dark matter where the DM scatters off the bath particles through $s$-wave and $p$-wave scattering, and compute the free-streaming horizon in terms of microscopic model parameters, in the presence of an EMDE with temperature-dependent entropy injection.